OpenStudy (anonymous):

solve the equations y^2 = 10y

OpenStudy (anonymous):

\[y^2-10y=0\] \[y(y-10)=0\]

OpenStudy (anonymous):

btw its not equations its only one equation

OpenStudy (lilg132):

y = ?

OpenStudy (anonymous):

well i have y = 0, 10 y = 0, -10 y = 10

OpenStudy (anonymous):

y(y-10)=0 one of them should be equal 0

OpenStudy (amistre64):

y.y = 10.y